OpenStudy (brenelle):

|v+8|-5=2 can some one help with that please

OpenStudy (sshayer):

\[\left| v-8 \right|-5=2\] \[\left| v-8 \right|=2+5\] \[v-8=\pm7,v=8 \pm 7,\] find the two values of v.

OpenStudy (brenelle):

Ok thank you

OpenStudy (brenelle):

The question was v+8 -5=2

OpenStudy (sshayer):

if no absolute value then only one solution.

OpenStudy (zzr0ck3r):

@NvidiaIntely

OpenStudy (loveneenah):

It is V=18/5 or if u need it in decimal form it will be 3.6>. hH